let f(x) be aquaratic eqn which is positive for all real x.then


f(x)+f'(x)+f"(x)is 


a)>0


b)>=0


c)=0


d)<0

Grade:12

1 Answers

Askiitians Expert Sahil Arora - IITD
13 years ago

Dear Rohit

f(x)=ax2 +bx+c

since it is positive for all x

a>0 and D<0==>b2 - 4ac<0

f(x)+f'(x)+f"(x)=ax2 +bx+c+ 2ax+b+2a=ax2 +(2a+b)x+2a+b+c

D=(2a+b)2 -4a(2a+b+c)=b2-4ac-4a2 <0 and a>0

==>always positive

Please feel free to post as many doubts on our discussion forum as you can. If you find any question 
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best rohit !!!




Regards,

Askiitians Experts

Sahil Arora